Based on the following reaction, calculate the vol
Answer #1

4) As we know the ideal gas equation as

PV=nRT

n= number of moles =\frac{w}{MW}

w= given mass

MW= molar mass

R= gas constant

T = temperature in Kelvin

Now here we have the reaction as

5O_3_{(g)}+2KOH\rightarrow 2KO_3+5O_2_{(g)}+H_2O_{(l)}

Here 2 mole of KO_3_{(s)} are produced from 5 mole of O_3_{(g)}

1 mole = molar mass in gm

2 mole of KO_3_{(s)} = 2 x 87.1 = 174.2 gm

5 mole of O_3_{(g)} = 5 x 48 = 240 gm

That means 174.2 gm of KO_3_{(s)} is produced from 240 gm of O_3_{(g)}

then 7.75 gm of KO_3_{(s)} will be produced from the amount of O_3_{(g)} as

\\=\frac{240}{174.2}\times 7.75\\ =10.67gm

Now the moles of O_3_{(g)} found as

\\=\frac{w}{MW}\\ =\frac{10.67}{48}\\ =0.22 mol

Again we know

R=62.36LTorrK^{-1}Mol^{-1}

P = 110 torr

n = 0.22 moles

T = 31 0 C = 31+273 =304 K

Now apply

PV=nRT

\\110torr\times V=0.22mol\times 62.36 LtorrK^{-1}mol^{-1}\times 304K\\ 110\times V=4170 L\\ V=\frac{4170.63}{110}L\\ V=37.91 L

Hence the volume of O_3_{(g)} is 37.91 L

Answer.